Calculate the energy required to ionise a hydrogen atom when the electron is residing in the third orbit ?

Grade:12th pass

1 Answers

Arun
5 years ago

The expression of energy is given by,

 

Where,

Z = atomic number of the atom

n = principal quantum number

For ionization from n1 = 5 to ,

Therefore ΔE = E2- E1 = - 21.8 X10-19 (1/n22-1/n12)

= 21.8 X10-19(1/n22-1/n12)

= 21.8 X10-19 (1/52-1/∞)

= 8.72 x 10-20 J

For ionization from 1st orbit, n1= 1, 

Therefore ΔE’ = 21.8x10-19(1/12-1/∞)

= 21.8x10-19 J

Now ΔE’/ ΔE = 21.8x10-19 / 8.72x10-20 = 25

Thus the energy required to remove electron from 1st orbit is 25 times than the required to electron from 5th orbit.